# Vulnerability of Text-Matching in ML/AI Conference Reviewer Assignments to Collusions

### v1

@@ 5 Results
In this section, we present the results of our experiments. We explain the experimental setup in Section 5.1. In Section 5.2, we simulate realistic human-in-the-loop attack scenarios and evaluate the attack effectiveness on 25 randomly selected (paper, reviewer) pairs of colluders. Then, we conduct larger-scale experiments using automatic abstract modifications and investigate the attack effectiveness even when colluding reviewers have low natural rankings (Section 5.3). In Section 5.4, we perform ablation studies to better understand the attack budgets. We also discuss scenarios where the attack success rates can be reduced in Sections 5.5 and 5.6. Furthermore, we discuss the potential usage of publicly released reviewer pool data for attack refinement in Section 5.7. Finally, in Section 5.8 we discuss a human subject experiment testing for the suspiciousness and detectability of the modified abstracts. All adversarial abstracts generated for all experiments can be found on GitHub (https://github.com/passionfruit03/reviewer_assignments_vulnerability/).
@@ 5.1 Experiment Setup
Using the OpenReview API (https://api2.openreview.net), we download all accepted papers at the NeurIPS 2023 venue. We download the names of all reviewers at NeurIPS 2023 (https://neurips.cc/Conferences/2023/ProgramCommittee) and search for OpenReview profiles that match the names of each reviewer. When curating the reviewer pool, we discard some reviewers if (1) there are multiple profiles that match the name of a reviewer or if (2) the reviewer has no public publications on their OpenReview profile. In this manner, we obtain 3,218 papers and 7,900 reviewers for the experiments. Following the same procedure, we also curate a NeurIPS 2022 dataset with 2,671 papers and 6,634 reviewers. We only download and use the paper metadata where Creative Commons Public Domain Dedication (CC0 1.0) apply (see Definition “Metadata” and Section “Metadata Dedication” in https://openreview.net/legal/terms).
@@
We are interested in the similarity of the colluding reviewer to the colluding paper compared to all other reviewers at the conference. Our evaluations are based on the competition rankings of reviewer similarities, where reviewers that have equal similarities receive the same ranking number, and then a gap is left in the ranking numbers (e.g. “1,2,2,4”). For any paper $p$ and reviewer $r$ at a conference, we define the natural ranking of ($p$, $r$) as the competition ranking, of $r$ when all reviewers at the conference are ranked by their similarity scores with the paper $p$. We use the term manipulated ranking for the colluding reviewer’s competition ranking by similarity to the colluding paper as a result of their manipulations.
@@
From the curated NeurIPS 2023 papers and reviewers, for each paper $p$ we find reviewers (there may be multiple since competition ranking is used) with natural rankings of 101. If no reviewer ranked 101st for a paper due to ties, we find the reviewers who have the next rank after 100. From all (paper $p$, reviewer $r$) pairs collected this way, we randomly sample without replacement 100 $(p,r){}$ pairs to act as colluders running the proposed attack algorithm. We evaluate the attack efficacy on these colluding paper-reviewer pairs. OpenReview sets the similarity score to be 00 if reviewer $r$ has a natural ranking greater than 100 for paper $p$, so it would be highly unlikely for $r$ to be assigned to review $p$ naturally. Following the same procedure, we also sample $(p,r){}$ pairs with natural rankings of 501 and 1001 for evaluation.
@@
We evaluate the effectiveness of our attack by measuring the top-1, top-3, and top-5 success rates in NeurIPS 2023, where a top-N success rate is defined as the fraction of times the proposed attack successfully increases the colluding reviewer’s manipulated ranking to be top-N for the colluding paper. We study these success rates because most papers are assigned to 3–6 reviewers at conferences [Sha22].
@@
The colluding reviewer constructs $\mathcal{Q}_{r}^{\text{adv}}$ by selecting the most similar paper to $p$ from the default archive $\mathcal{Q}_{r}$ of up to 10 most-recent papers they have authored (Section 4.1). In addition, the colluding author modifies their adversarial abstract $a^{\text{adv}}_{p}$ to be more similar to $\mathcal{Q}_{r}^{\text{adv}}$ (Section 4.2). We investigate human-in-the-loop attack efficacy with early stopping in Section 5.2. In subsequent sections, we investigate fully automatic attack results and do not use early stopping.
@@
There are three hyperparameters $\bm{N},{\bm{M}},{\bm{K}}$ we use to define the attack budget. In IncludeThemes, $\bm{N}$ stands for the number of $a^{\text{adv}}_{p}$ versions created before selecting the most similar version. In InsertKeywords, ${\bm{M}}$ is the number of batches of keywords to insert, and ${\bm{K}}$ is the maximum number of keywords in each batch. We explore the attack success rates under different combinations of $\bm{N},{\bm{M}},{\bm{K}}$ with the NeurIPS 2022 dataset and select the highest performing combination $\bm{N}=5$, ${\bm{M}}=2$, ${\bm{K}}=5$. In Section 5.4, we also present an investigation into different choices of hyperparameters $\bm{N},{\bm{M}},{\bm{K}}$ on the NeurIPS 2023 dataset.
@@
5.2 Human-in-the-loop Mode Attack Success Rate in NeurIPS 2023
We perform 25 attacks by constructing the $\mathcal{Q}_{r}^{\text{adv}}$ and modifying $a^{\text{adv}}_{p}$ in the more realistic human-in-the-loop mode, then we evaluate the attack success rates. We randomly sample (p,r) pairs with natural rankings of 101, and we keep the first 25 samples with paper topics we are familiar enough with to judge the coherence and consistency of the modified abstracts. Firstly, we use the method described in Section 4.1 to construct $\mathcal{Q}_{r}^{\text{adv}}$. For the abstract modifications, we simulate what colluding authors would do when abstracts are modified with human involvement. The human-in-the-loop implementations of IncludeThemes and InsertKeywords are described in Algorithm 1 and Algorithm 2, respectively. In addition, we do early stopping checks to prevent abstracts from being modified more than necessary. The exact early stopping heuristics used can be found in Section 4.2.
@@ 5.2 Human-in-the-loop Mode Attack Success Rate in NeurIPS 2023
In addition to increasing $\bm{N}$, we allow up to 3 sentences about the colluding reviewer’s archive $\mathcal{Q}_{r}$ to be added to the abstract during IncludeThemes, instead of the 1 sentence constraint in the fully automatic mode. This is a realistic change because coherence and consistency are enforced manually in this mode, and additional sentences often improve abstract flow by allowing better transition sentences. As for FindKeywords and InsertKeywords, we choose the default values of ${\bm{K}}=5$ and ${\bm{M}}=2$. In fact, because of the early stopping, most abstracts have less than 10 keywords added.
@@
Natural Ranking Top-1 Top-3 Top-5 Mean
95% CI
@@
5.3 Fully Automatic Mode Attack Success Rates in NeurIPS 2023
In the remainder of this paper, we investigate the success rates of our attack under fully automatic mode abstract manipulation without early stopping. The fully automatic implementation of IncludeThemes and InsertKeywords can be found in Algorithm 1 and Algorithm 2, respectively.
We randomly sample without replacement 100 (paper $p$, reviewer $r$) pairs each for three scenarios, where $r$ has a natural ranking of 101st, 501st, or 1001st for $p$, respectively. We simulate collusion between the paper and the reviewer and run the proposed attack procedure for each pair, and then we evaluate the attack success rates among the 100 samples for each scenario. Again, we use the method described in Section 4.1 to construct $\mathcal{Q}_{r}^{\text{adv}}$ first. The abstract modifications are made under the fully automatic mode.
Table 2: Attack success rates in fully automatic mode for colluding reviewers with natural rankings of 101, 501, and 1001.
@@ 5.3 Fully Automatic Mode Attack Success Rates in NeurIPS 2023
Natural Rankings Top-1 Top-3 Top-5 Mean
95% CI
@@
We enumerate the success rates of the attack in Table 2. We find that the success rates are generally high. For example, when the natural ranking is 101, the proposed attack procedure leads to a top-5 attack success rate of 88%. Even when the natural ranking is 1001, the proposed attack procedure can successfully increase the manipulated ranking to be 1st in 48% of the cases (top-1 success rate). These results highlight the potential that colluding reviewers who are not working directly in the same area can still collude to successfully manipulate reviewer assignments.
5.4 Attack Budgets
To help us understand the amount of change that is needed to be made to the abstract in order to for attacks to succeed, we explore the attack success rates under different attack budgets on $(p,r)$ samples from the NeurIPS 2023 dataset. We study the effectiveness of each of the two abstract modification operations — IncludeThemes and InsertKeywords in increasing the similarity $s((t_{p},a^{\text{adv}}_{p}),\mathcal{Q}_{r}^{\text{adv}})$ between the adversarial abstracts and the curated archive $\mathcal{Q}_{r}^{\text{adv}}$. We investigate the effect of the choice of hyperparameters $\bm{N}$, ${\bm{M}}$, and ${\bm{K}}$ representing the attack budgets used in IncludeThemes and InsertKeywords (Section 4.2). We report the fully automatic mode attack success rates of different values of $\bm{N}$, ${\bm{M}}$ and ${\bm{K}}$ on 50 samples with natural rankings of 101. First, in Figure 4, we report the attack success rates for a range of values for $\bm{N}$ – the number of rounds for which IncludeThemes is performed – without performing the InsertKeywords operation (${\bm{M}}=0$, ${\bm{K}}=0$). We show that attack success rates generally increase with $\bm{N}$ in Figure 4. This is expected because language model embeddings can be sensitive to paraphrasing, so taking the most-similar attempt amongst stochastic outputs helps improve attack success rates. We also note that there are signs of diminishing returns as $\bm{N}$ increases.
Next, we report the attack success rates with different values of parameters ${\bm{M}}$ and ${\bm{K}}$ for the InsertKeywords operation and the FindKeywords routine, without performing the IncludeThemes operation ($\bm{N}=0$). In Figure 5, we find that attack success rates increase with the total number of keywords incorporated ${\bm{M}}\times{\bm{K}}$. This trend can be seen through the color gradients from bottom left to top right in the subfigures of Figure 5. In addition, for a fixed number of keywords inserted overall, we notice that inserting smaller batches of keywords over more InsertKeywords iterations is helpful when compared to inserting a larger batch of keywords over less InsertKeywords iterations. This can be seen by comparing the reflected squares between upper-left and bottom-right of the subfigures in Figure 5. This result aligns with our intuition about finding keywords in smaller batches and inserting them over more iterations can improve the greedy search for a new batch of keywords by taking into account the new changes to $a^{\text{adv}}_{p}{}$ from inserting the previous batches of keywords.
Figure 4: Attack success rates generally increase with $\bm{N}$ when the most similar attempt is kept out of $\bm{N}$ versions of $a^{\text{adv}}_{p}$ in IncludeThemes. The band represents standard error of the mean.
(a) Top-1 success rates
(b) Top-3 success rates
(c) Top-5 success rates
5.5 Lower Limits on Reviewer’s Archive Length
@@
What is the effect of simply curating the adversarial archive $\mathcal{Q}_{r}^{\text{adv}}$ as described in Section 4.1, without any manipulation of abstractions by authors ($a^{\text{adv}}_{p}{}=a_{p}$)? For colluding $(p,r){}$ pairs with natural rankings of 101, we find that 33% of the samples have manipulated rankings being within the top-3 from adversarial reviewer archive curation alone—see Table 3. In addition, the mean manipulated ranking is 24.59, which is much improved compared to the natural ranking of 101. This is in fact a significant vulnerability, and reviewer archive curation alone is a serious threat to automated reviewer assignments.
@@
Natural Ranking Top-1 Top-3 Top-5 Mean
95% CI
@@
[18.01, 31.17]
A possible defense to our attack could be imposing a lower limit on the number of publications each reviewer has to keep in their archive. To investigate such a defense, we randomly sample without replacement 100 $(p,r)$ pairs to act as colluders with natural rankings of 101 and where the reviewer $r$ has at least ten publications. For the curation of $\mathcal{Q}_{r}^{\text{adv}}$, we consider four different scenarios where the reviewers keep the 1, 2, 5, or 10 most-similar publications in $\mathcal{Q}_{r}^{\text{adv}}$. Afterwards, we run automatic abstract modification and evaluate the success rates for each scenario. Figure 6 shows that attack success rates decreases with $\lvert\mathcal{Q}_{r}^{\text{adv}}\rvert$, meaning that imposing a high lower limit on the reviewer’ archive lengths can effectively decrease the proposed attack’s success rates. However, there is a trade-off here, since honest reviewers may actually want to update their profiles to reflect their most current research interests.
Figure 6: Attack success rates when the colluding reviewers have to keep 1, 2, 5, 10 papers in the adversarial archive $\lvert\mathcal{Q}_{r}^{\text{adv}}\rvert$. Success rates drop when colluding reviewers must keep more papers in their archive. The shaded bands represent standard errors of the mean.
@@ 5.6 Maximum versus Average Similarity
Aggregation Method Top-1 Top-3 Top-5 Mean
95% CI
@@
In the case where the colluding reviewer $r$ is not allowed to make changes to their archive (that is, $\mathcal{Q}_{r}^{\text{adv}}=\mathcal{Q}_{r}$), we hypothesize that our attack would be more effective against the maximum aggregation method since abstract modifications can target just one paper in $\mathcal{Q}_{r}$. We randomly sample without replacement 100 (paper $p$, reviewer $r$) pairs with natural rankings of 101 under each aggregation method. Then, for each pair, we run the automatic abstract modification attack procedure on the paper $p$ to increase the similarity to reviewer $r$ (no adversarial archive curation). We evaluate the attack success rates amongst the 100 samples for each scenario and enumerate our results in Table 4. We indeed find that our proposed attack is more successful under maximum aggregation. This result suggests that conferences that choose to use the more popular method of maximum aggregation method are generally more susceptible to reviewer assignment manipulation using our proposed attack.

### v2

@@ 5 Results
In this section, we present the results of our experiments. We explain the experimental setup in Section 5.1. In Section 5.2, we simulate realistic human-in-the-loop attack scenarios and evaluate the attack effectiveness on 25 randomly selected (paper, reviewer) pairs of colluders. Then, we conduct larger-scale experiments using automatic abstract modifications and investigate the attack effectiveness even when colluding reviewers have low natural rankings (Section 5.3).
We also discuss scenarios where the attack success rates can be reduced in Sections 5.4 and 5.5. Furthermore, we discuss the potential usage of publicly released reviewer pool data for attack refinement in Section 5.6. Finally, in Section 5.7 we discuss a human subject experiment testing for the suspiciousness and detectability of the modified abstracts. All adversarial abstracts generated for all experiments are publicly downloadable from Zenodo at https://doi.org/10.5281/zenodo.15588237.
@@ 5.1 Experiment Setup
We download all accepted papers at the NeurIPS 2023 venue using the OpenReview API (https://api2.openreview.net). We download the names of all reviewers at NeurIPS 2023 (https://neurips.cc/Conferences/2023/ProgramCommittee) and search for OpenReview profiles that match the names of each reviewer. When curating the reviewer pool, we discard some reviewers if (1) there are multiple profiles that match the name of a reviewer or if (2) the reviewer has no public publications on their OpenReview profile. In this manner, we obtain 3,218 papers and 7,900 reviewers for the experiments. Following the same procedure, we also curate a NeurIPS 2022 dataset with 2,671 papers and 6,634 reviewers. We only download and use the paper metadata where Creative Commons Public Domain Dedication (CC0 1.0) apply (https://openreview.net/legal/terms).
@@
Our evaluations are based on the competition rankings of reviewer similarities, where reviewers that have equal similarities receive the same ranking number, and then a gap is left in the ranking numbers (e.g. “1,2,2,4”). For any paper $p$ and reviewer $r$ at a conference, we define the natural ranking of ($p$, $r$) as the competition ranking of $r$ when all reviewers at the conference are ranked by their similarity scores with the paper $p$, without manipulations. We use the term manipulated ranking for the colluding reviewer’s competition ranking by similarity to the colluding paper as a result of their manipulations. Appendix C relates natural rankings to absolute similarity.
@@
From the curated NeurIPS 2023 papers and reviewers, for each paper $p$, we find reviewers (there may be multiple since competition ranking is used) with natural rankings of 101. If no reviewer ranked 101st for a paper due to ties, we find the reviewers who have the next rank after 100. From all (paper $p$, reviewer $r$) pairs collected this way, we randomly sample without replacement a number of (p,r)(p,r){} pairs to act as colluders running the proposed attack algorithm. We evaluate the attack efficacy on these colluding paper-reviewer pairs. OpenReview sets the similarity score to be 0 if reviewer $r$ has a natural ranking greater than 100 for paper $p$, so it would be highly unlikely for $r$ to be assigned to review $p$ naturally. Similarily, we also sample (p,r)(p,r){} pairs with natural rankings of 20, 501 and 1001 for evaluation. We specify exact sample sizes in each section.
@@
We evaluate the effectiveness of our attack by measuring the top-1, top-3, and top-5 success rates in NeurIPS 2023, where a top-N success rate is defined as the fraction of times the attack successfully increases the colluding reviewer’s manipulated ranking to be top-N for the colluding paper. We study these success rates because most papers are assigned to 3–6 reviewers at conferences [Sha22].
@@
The colluding reviewer constructs $Q_{r}^{\text{adv}}$ by selecting the most similar paper to $p$ from the default archive $Q_{r}$ of up to 10 most-recent papers they have authored (Section 4.1). In addition, the colluding author modifies their adversarial abstract $a^{\text{adv}}_{p}$ to be more similar to $Q_{r}^{\text{adv}}$ (Section 4.2). We investigate human-in-the-loop attack efficacy with early stopping in Section 5.2. In subsequent sections, we investigate fully automatic attack results and do not use early stopping. The fully automatic mode does not require human supervision, which makes large-scale evaluations and further ablations feasible.
@@
There are three hyperparameters $\bm{N},{\bm{M}},{\bm{K}}$ we use to define the attack budget. In IncludeThemes, $\bm{N}$ stands for the number of $a^{\text{adv}}_{p}$ versions created before selecting the most similar version. In InsertKeywords, $\bm{M}$} is the number of batches of keywords to insert, and $\bm{K}$} is the maximum number of keywords in each batch. We explore the attack success rates under different $\bm{N},{\bm{M}},{\bm{K}}$ with the NeurIPS 2022 dataset and select the highest performing combination $\bm{N}=5$, ${\bm{M}}=2$, ${\bm{K}}=5$. In Appendix B, we present an investigation into different choices of hyperparameters $\bm{N},{\bm{M}},{\bm{K}}$ on the NeurIPS 2023 dataset.
@@
5.2 Human-in-the-loop Mode Attack Results
We perform 25 attacks by constructing the $Q_{r}^{\text{adv}}$ and modifying $a^{\text{adv}}_{p}$ in the more realistic human-in-the-loop mode. We randomly sample (p,r) pairs with natural rankings of 101, and we keep the first 25 samples with paper topics we are familiar enough with to judge the coherence and consistency of the modified abstracts. Firstly, we use the method described in Section 4.1 to construct $Q_{r}^{\text{adv}}$. For the abstract modifications, we simulate what colluding authors would do when abstracts are modified with human involvement. The human-in-the-loop implementations of IncludeThemes and InsertKeywords are described in Appendix A Algorithm 1 and Algorithm 2, respectively. In addition, we do early stopping checks to prevent abstracts from being modified more than necessary. The exact early stopping heuristics used can be found in Section 4.2 and Appendix A.
@@ 5.2 Human-in-the-loop Mode Attack Results
In addition to increasing $\bm{N}$, we allow up to 3 sentences about the colluding reviewer’s archive $Q_{r}$ to be added to the abstract during IncludeThemes, instead of the 1 sentence constraint in the fully automatic mode. This is a realistic change because coherence and consistency are enforced manually here, and additional sentences often improve abstract flow by allowing better transition sentences. As for FindKeywords and InsertKeywords, we choose the default values of ${\bm{K}}=5$ and ${\bm{M}}=2$. In fact, because of the early stopping, most abstracts have less than 10 keywords added.
@@
Natural Top-1 Top-3 Top-5 Mean
Rankings 95% CI
@@
5.3 Fully Automatic Mode Attack Results
In the remainder of this paper, we investigate the success rates of our attack under fully automatic mode abstract manipulation without early stopping. The fully automatic implementation of IncludeThemes and InsertKeywords can be found in Algorithm 1 and Algorithm 2, respectively. Unlike the human-in-the-loop results in Section 5.2, the fully automatic mode involves no human supervision of the LLM outputs. Our motivation for using the fully automatic mode is to enable large-scale, reproducible evaluations.
We randomly sample without replacement 300 (paper $p$, reviewer $r$) pairs where $r$ has a natural ranking of 101 for $p$. For the natural ranking of 20, 501 and 1001, we sample 100 $(p,r)$ pairs. (The larger sample size for the 101st scenario is due to a request from a reviewer of this paper.) Afterwards, we perform the attack procedure for each $(p,r)$ pair by curating the adversarial archive $Q_{r}^{\text{adv}}$ (Section 4.1) and then performing fully automatic mode abstract modifications (Section 4.2).
We enumerate the attack success rates in Table 2. We find that the success rates are generally high. When the natural ranking is 20, the attack success rates are the highest since the colluding $(p,r)$ pair is naturally highly similar. When the natural ranking is 101, the proposed attack procedure leads to a top-5 attack success rate of 93%. Even when the natural ranking is 1001, the top-1 attack success rate is 48%. These results highlight that colluding $(p,r)$ pairs can successfully manipulate reviewer assignments, whether they work on similar topics or not. Appendix D further explore the relationships between success rates and absolute similarity.
Table 2: Attack success rates in fully automatic mode for colluding reviewers with natural rankings of 20, 101, 501, and 1001.
@@ 5.3 Fully Automatic Mode Attack Results
Natural Top-1 Top-3 Top-5 Mean
Rankings 95% CI
@@
5.4 Lower Limits on Reviewer Archive Length
@@
What is the effect of simply curating the adversarial archive $Q_{r}^{\text{adv}}$ as described in Section 4.1, without any manipulation of abstracts by authors ($a^{\text{adv}}_{p}{}=a_{p}$)? For colluding (p,r)(p,r){} pairs with natural rankings of 101, we find that 30% of the samples have manipulated rankings being within the top-3 from adversarial reviewer archive curation alone—see Table 3. In addition, the mean manipulated ranking is 24.59, which is much improved compared to the natural ranking of 101. This is in fact a significant vulnerability, and reviewer archive curation alone is a serious threat to automated reviewer assignments.
A possible defense to our attack could be imposing a lower limit on the number of publications each reviewer has to keep in their archive. To investigate such a defense, we randomly sample without replacement 100 $(p,r)$ pairs to act as colluders with natural rankings of 101 and where the reviewer $r$ has at least ten publications. For the curation of $Q_{r}^{\text{adv}}$, we consider four different scenarios where the reviewers keep the 1, 2, 5, or 10 most-similar publications in $Q_{r}^{\text{adv}}$. Afterwards, we run automatic abstract modification and evaluate the success rates. Figure 4 shows that attack success rates decreases with $\lvert Q_{r}^{\text{adv}}\rvert$, meaning that imposing a high lower limit on the reviewer’ archive lengths can effectively decrease the proposed attack’s success rates. However, there is a trade-off here, since honest reviewers may actually want to update their profiles to reflect their most current research interests.
@@
Natural Top-1 Top-3 Top-5 Mean
Rankings 95% CI
@@
[19.16, 26.45]
Figure 4: Attack success rates when the colluding reviewers have to keep 1, 2, 5, 10 papers in the adversarial archive $\lvert Q_{r}^{\text{adv}}\rvert$. Success rates drop when colluding reviewers must keep more papers in their archive. The shaded bands represent standard errors of the mean.
@@ 5.5 Maximum versus Average Similarity
Aggregati- Top-1 Top-3 Top-5 Mean
on Method 95% CI
@@
In the case where the colluding reviewer $r$ is not allowed to make changes to their archive (that is, $Q_{r}^{\text{adv}}=Q_{r}$), we hypothesize that our attack would be more effective against the maximum aggregation method since abstract modifications can target just one paper in $Q_{r}$. We randomly sample without replacement 100 (paper $p$, reviewer $r$) pairs with natural rankings of 101 under each aggregation method. Then, for each pair, we run the automatic abstract modification attack procedure on the paper $p$ to increase the similarity to reviewer $r$ (no adversarial archive curation). We evaluate the attack success rates amongst the 100 samples for each scenario and enumerate our results in Table 4. We indeed find that our proposed attack is more successful under maximum aggregation. This result suggests that conferences that choose to use the more popular maximum aggregation method are generally more susceptible to our proposed reviewer assignment attack.